A construction crew is lengthening a road that originally measured 52 miles. The crew is adding one mile to the road each day. The length, L(in miles), after D days of construction is given by the following function L(d)=52+b. What is the length of the road after 36 days
Answer by ikleyn(53748)

.
A construction crew is lengthening a road that originally measured 52 miles. The crew is adding one mile
to the road each day. The length, L(in miles), after D days of construction is given by the following
function L(d)=52+b. What is the length of the road after 36 days
~~~~~~~~~~~~~~~~~


            The function in your post is written incorrectly.

            A correct writing is   L(D) = 52 + D,   where  D  is the number of working days.


To get the answer, simply substitute 36 instead of D into the formula 

    L(36) = 52 + 36 = 88 miles.    ANSWER

Solved.

--------------------

This problem is extremely simple,  but is written with a lot of pathos.

When there are so much pathos,  it imposes some obligations:
at least,  the text  (including the formulas)  should be correct.